OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

Can you do the parts one at a time please.

OpenStudy (anonymous):

p1 you read my mind :) one at time y-y1=m(x-x1) is your formula for point slope form -6-y1=4(3-x1) i think that is all for p1

OpenStudy (anonymous):

Okay I understand that part.

OpenStudy (anonymous):

okey then now p2

OpenStudy (anonymous):

p2 take -6-y1=4(3-x1) and solve for y basicaly that is what is meant by slope intercecpt form y=mx+b -6-y1=4(3-x1) sfy

OpenStudy (anonymous):

Minus the 6 from each side?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

-6-y1=4(3-x1) sfy actuall it is add the 6 since it is - NOTE you can change the name of variable at anytime so lets call y1 y and x1 x -y=-1*(4(3-x) +6) mult by -1 y=-4(3-x)-6 distribute -1 y=-12+4x-6 distubute 4 y=-18+4x add like terms 12 and -6 y=4x-18 rearrange THat is part 2

OpenStudy (anonymous):

part 3 equation is like ax+by=c y=4x-18 18= 4x- y < c = ax- y I think that is right